### v1

@@ 4 Algebraic versions of the combinatorial BFS and CCS
  • Finding the connected $\textnormal{\bf Finding the connected component}\ (G,s\!\in\!V):\mathcal{C};$
@@
Let us show that iterations (6) give us the graph traversal algorithm, which is equivalent to the combinatorial BFS, and that iterations (7) give us the graph traversal algorithm, which is equivalent to the combinatorial CCS.
@@ 5 Jacobi iterations as implementation
Considering BFS, for starting vertex $s$ and $i\!\in\!\mathcal{F}^{(k+1)}$, let $C(s,i)$ be the set of simple chains, which connect $s$ with $i$. All of the chains, if there is any, have the length $k$.
@@
$x_{i}^{(k+1)}\!=\!\biggl{(}\sum\limits_{j=1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}(-d)=\biggl{(}\sum\limits_{(i,j)\in E}\biggl{(}\sum\limits_{c\in C(s,j)}(-1)^{k}d^{k+1}\biggl{)}\biggl{)}(-d)=$
$=\biggl{(}\sum\limits_{(i,j)\in E}\biggl{(}\sum\limits_{c\in C(s,j)}(-1)^{k+1}d^{k+2}\biggl{)}\biggl{)}=\sum\limits_{c\in C(s,i)}(-1)^{k+1}d^{k+2}\neq 0.$
So, the condition (5) is met for $i\!\in\mathcal{F}^{(k+1)}$. Let us show that $x_{i}^{(k+1)}\!=\!0$ if $i\!\not\in\mathcal{F}^{(k+1)}$ and the vertex $i$ was not visted at previous $k$ iterations of the combinatorial BFS. This means that among the vertices $j$ adjacent to $i$ there are no ones for which there have been built simple chains of length $k$ connecting $s$ and $j$ in the course of the previous $k$ iterations. Since Lemma 1 is true for $k$, this means that for all such $j$ we have $x_{j}^{(k)}\!=\!0$. Therefore,
@@
Thus, by Lemma 1 and by (5), the iterations (6) produce the same frontiers $\mathcal{F}^{(k)}$, $k\!=\!1,2,\ldots$, as iterations of the combinatorial BFS produce. This means that we obtain the same graph traversals during execution of the iterations. Theorem 1 follows from that.
@@ 6 Gauss-Seidel iterations as implementation
6.1 Computing components of the state vector through the chains outgoing from the current frontier
@@
Considering the $(k\!+\!1)$-th iteration of the combinatorial CCS and iteration of the form (7) with the same number, we will use the following notation. Unlike to the previous section, we denote as $C(\tilde{s},i)$ the set of simple chains connecting the vertices $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ with vertices $i\!\in\!\mathcal{F}^{(k+1)}$. Along edges of these chains transitions are made at $(k\!+\!1)$-th iteration of the combinatorial CCS. Let $c\!+\!(j,i)$ denotes the chain obtained from $c$ by connecting its last vertex $j$ to vertex $i$ by the edge $(j,i)\!\in\!E$.
All computations in iterations (6) and (7) can be decomposed into computations that implemented in accordance with traversals of individual chains (not necessary simple) connecting the starting vertex and the vertex $i$, for which the value $x_{i}^{(k+1)}$ is computing.
@@
We shall consider only simple chains since it will be enough for the proofs of the followed lemma and theorem. Indeed, when implemeting computations of the form (6) or (7), the computations for some set of walks from starting vertex to other vertices are constitute the set of all individual operations with machine numbers implemented in the course of one iteration.
@@
Chains that can be traversed in the course of one iteration of combinatorial CCS, that is, chains $c\!=\!\{i_{0},i_{1},\ldots,i_{\ell(c)}\}\!\in\!C(\tilde{s},i)$, $i_{0}\!=\!\tilde{s}$, are chains of the two types. A chain $c\!\in\!C(\tilde{s},i)$ is a chain of type (I) if $i_{0}\!<\!i_{1}$, and the chain type is (II) if $i_{0}\!>\!i_{1}$. For both type chains, maybe except the first edge $(i_{0},\!i_{1})$, we have $i_{j-1}\!<\!i_{j}$ for $j\!=\!\overline{1,\ell(c)}$. Transition along the edge $(i_{0},i_{1})$ is implemented at step 3 of the combinatorial CCS and transitions along subsequent edges of the chain are performed at the step 4.
@@
For the vertices $\tilde{s},i\!\in\!V$, let $\ell(\tilde{s},i)\!=\!\max\{\ell(c)\,|\,c\in C(\tilde{s},i)\}$. $\ell(\tilde{s},i)$ is the maximum length of a simple chain outgoing from $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ along which $i\!\in\!\mathcal{F}^{(k+1)}$ is reached at the $(k\!+\!1)$-th iteration.
@@
Let $i_{0}\!=\!\tilde{s}$, $i_{\ell(c)}\!=\!i$ in the chain $c\!=\!\{i_{0},\ldots,i_{\ell(c)}\}$ and $v_{\tilde{s}}$ is the value that is transmitted through the chain $c$ that connects the vertex $\tilde{s}$ and the vertex $i$. As a result of this transmission of $v_{\tilde{s}}$, by traversing the chain $c$, we obtain the contribution $x_{i,c}^{(k+1)}$ into the whole value of $x_{i}^{(k+1)}$ by this chain. There must be a chain through which the non-zero value $v_{\tilde{s}}$ is transmitted to $i$ to have $x_{i}^{(k+1)}\!=\!0$.
@@
So, for $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ and $c\!\in\!C(\tilde{s},i)$ we define $x_{i,c}^{(k+1)}$ as the contribution of the value $v_{\tilde{s}}$, which transmitted along the chain $c\!\in\!C(\tilde{s},i)$ from $\tilde{s}$ to $i$:
@@
If $C(\tilde{s},i)\!=\!\varnothing$ then $x_{i(\tilde{s})}^{(k+1)}\!=\!0$ since there are no chains, through which the contributution $v_{\tilde{s}}$ can be transmitted to $i$ from $\tilde{s}$. The value $x_{i(\tilde{s})}^{(k+1)}$ is the sum of all contributions to the value of $x_{i}^{(k+1)}$, which are transmitted through all chains in $C(\tilde{s},i)$ at the course of the $(k\!+\!1)$-th iteration (7).
Note that during implementation of the algebraic CCS, reaching vertices from $\mathcal{F}^{(k+1)}$ can be done not only along simple chains. Traversal of a chain of the type (II) can lead to the fact that a correct chain that otgoes from some vertex $j$ already reached at the course of an iteration (7), at the same iteration will repeatedly pass through the vertex $\tilde{s}$ from which $j$ was reached. The value obtained along traversals of a such loops will be added to the initial value of $x_{\tilde{s}}^{(k)}$, further transmitted along the chains of $C(\tilde{s},i)$. That is, it will be included in $v_{\tilde{s}}$ from (8) while $\tilde{s}$ is revisited. This happens when $\tilde{s}\!>\!j$.
Consider the examples we have on Fig. 1. On the pic. 1. a) and pic. 1. b) of the Fig. 1, the values of $x_{i}^{(k+1)}$ for the two first stages of the first iteration of the algebraic CCS (steps 3 and 4) for the graph are shown. The starting vertex is vertex $2$. Transmitting from the loop $c_{1}\!=\!\{2,1,2\}$, the value $v_{s}\!=\!d$ contributes $-d^{4}$ to $x_{3}^{(1)}$ through the whole walk $\{2,1,2,3\}$, which is traversed at the $1$-st iteration. Adding the contribution of the simple chain $c_{2}\!=\!\{1,3\}$, which constitutes of the one edge $(1,3)$, we obtain the final value of $x_{3}^{(1)}$ for the 1-st iteration as sum of contributions of the walk and the chain:
$x_{3}^{(1)}=x_{3,2}^{(1)}=x_{3,c_{2}}^{(1)}=\bigl{(}x_{2}^{(1)}+x_{2,c_{1}}^{(1)}\bigr{)}\cdot(-d)=\bigl{(}d+d^{3}\bigr{)}\cdot(-d)=-d^{2}-d^{4},$
since $x_{2}^{(1)}\!=\!d$, $x_{2,c_{1}}^{(1)}\!=\!d^{3}$.
The pic. 2. of the Fig. 1 shows how do the contributions coming to the vertex 5 along two different correct chains, which traversed at one iteration of CSS, add together. These chains are $c_{1}\!=\!\{1,3,4,5\}$ and $c_{2}\!=\!\{1,2,5\}$. The starting vertex is vertex $1$.
@@ 6.2 Gauss-Seidel iterations as implementation
$x_{i}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}x_{i(\tilde{s})}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}\sum\limits_{c\in C(\tilde{s},i)}x_{i,c}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}\sum\limits_{c\in C(\tilde{s},i)}v_{\tilde{s}}\cdot(-d)^{\ell(c)}.$ (11)11( 11 )
@@
Let us show that Lemma 2 is true for the $1$-st iteration, that is, for $k\!=\!1$. The frontier at the $1$-st iteration consists of a single starting vertex $s$: $\mathcal{F}^{(0)}\!=\!\{s\}$. Let $i\!\in\!\mathcal{F}^{(1)}$ be such that $\ell(s,i)\!=\!1$. Then $C(s,i)$ consists of a single chain $c\!=\!\{s,i\}$ of length $1$, i.e. it consists of the one edge $(s,i)\!\in\!E$. Thus we have $x_{i,c}^{(1)}\!=\!-d^{2}$ by (8), since $v_{s}\!=\!d$ at the $1$-st iteration (7).
For the $1$-st iteration (7), we have $x_{j}^{(0)}\!=\!0$ for all $j\!>\!i$, $j\!\neq\!s$. At the $1$-st iteration, we also have $x_{j}^{(1)}\!=\!0$ for $j\!<\!i$, $j\!\neq\!s$. If it is not holds, then $\ell(s,i)\!>\!1$, since there must be a chain $c\!\in\!C(s,j)$ with $\ell(c)\!>\!1$ through which the value $v_{s}\!=\!x_{s}^{(0)}$ is transmitted to $j$ by chain traversal algorithm.
Consequently, all the terms in (10) are zero except for $x_{s}^{(0)}$ or $x_{s}^{(1)}$, and we have
@@
where $\alpha\!=\!1$ if $(s,i)\!\in\!E$, and $\alpha\!=\!0$ otherwise. If $x_{j}^{(1)}\!\neq\!0$, then this means that either $j\!=\!s$ and $\alpha\!=\!1$ or, if $j\!\neq\!s$, we have $C(s,j)\!\neq\!\varnothing$. For all $c\!\in\!C(s,j)$, $\ell(c)\!\leq\!\ell(s,j)\!\leq\!l$ since $\ell(s,i)\!=\!l\!+\!1$. Therefore, by the induction assumption on $l$, we obtain:
@@
Since $x_{i,c}^{(1)}\!=\!x_{j,c^{\prime}}^{(1)}\cdot(-d)$ for $c^{\prime}\!\in\!C(s,j)$, $c\!=\!c^{\prime}\!+\!(j,i)$, $c\!\in\!C(s,i)$, then, substituting (13) into (12) and having $v_{s}\cdot(-d)\!=\!x_{i,c}^{(1)}$ for $c\!=\!\{s,i\}$, we obtain
@@
Consider $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!1$ for a vertex $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. For this case, all chains $c\!\in\!C(\tilde{s},i)$ are edges $(\tilde{s},i)\!\in\!E$, each of which, in (10), corresponds to a non-zero term of the form
@@
Suppose that it is true for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!\leq\!l$, $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. Let us show that it will also be true for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!l\!+\!1$.
@@
If $j\!<\!i$ and $x_{j}^{(k+1)}\!\neq\!0$, then $C(\tilde{s},j)\!\neq\!\varnothing$ by induction assumption on $k$. The length of each chain in $C(\tilde{s},i)$ is less than or equal to $l$, since otherwise $\ell(\tilde{s},i)\!>\!l\!+\!1$. Therefore, by induction on $l$, we have
@@
Consider the second term $S_{2}$ in (10):
@@
Adding $S_{1}$ and $S_{2}$, from (15) and (16) we obtain:
@@
This means that Lemma 2 is true at the $(k\!+\!1)$-th iteration for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!l\!+\!1$. Thus, by induction on $l$, Lemma 2 true for all vertices in $\mathcal{F}^{(k+1)}$, i.e., it is true for the $(k\!+\!1)$-th iteration. It means that, by the induction on $k$, it is true for an arbitrary $k$ ($k\!\leq\!n$).
@@
where $v_{\tilde{s}}\!=\!O(d^{\ell(s,\tilde{s})})\!\neq\!0$, and where $\ell(s,\tilde{s})$ is the maximum length of a simple chain connecting the starting vertex $s$ and the vertex $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. By definition (8), when traversing along each edge of a such chain, the value $v_{\tilde{s}}$, which is a sum of powers $d$ taken with some coefficients, is multiplied by $-d$. In this case, the same powers $d$ in the resulting values $x_{i,c}^{(k+1)}$ have the same sign, since they correspond to the same number of multiplications by $-d$ implemented by Chain traversal algorithm for chains of the same lengths. Therefore, if $C(\tilde{s},i)\!\neq\!\varnothing$, then $x_{i}^{(k+1)}\!\neq\!0$ and condition (5) is met for $i\!\in\!\mathcal{F}^{(k+1)}$ of the combinatorial CCS after the $(k\!+\!1)$-th iteration.
Let $i\!\in\!V\setminus\mathcal{C}^{(k+1)}$ for the combinatorial CCS, that is, vertex $i$ is not reached after the $(k\!+\!1)$-th iteration. This means that $C(\tilde{s},i)\!=\!\varnothing$ for every $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. By the proven Lemma 2, it follows from this that

### v2

@@ 4 Algebraic versions of the combinatorial BFS and CCS
  • Traversal of a connected $\textnormal{\bf Traversal of a connected component}\ (G,s\!\in\!V):\mathcal{C};$
@@
For the case when F is the transformation of the form (6), the algorithm is the variation of algebraic BFS. For F of the form (7), it gives algebraic CCS.
@@
Let us show that iterations (6) give us the graph traversal algorithm which is equivalent to the combinatorial BFS, and that iterations (7) give us the graph traversal algorithm which is equivalent to the combinatorial CCS.
@@ 5 Jacobi iterations as implementation
Considering BFS, for starting vertex $s$ and $i\!\in\!\mathcal{F}^{(k+1)}$, let $C(s,i)$ be the set of simple chains which connect $s$ with $i$. All of the chains, if there is any, have the length $k$, i.e., for all $c\!\in\!C(s,i)$ we have $\ell(c)\!=\!k$.
@@
$x_{i}^{(k+1)}\!=\!\biggl{(}\sum\limits_{j=1}^{n}a_{ij}x_{j}^{(k)}\biggr{)}(-d)=\biggl{(}\sum\limits_{(i,j)\in E}\biggl{(}\sum\limits_{c\in C(s,j)}(-1)^{\ell(c)}d^{\ell(c)+1}\biggl{)}\biggl{)}(-d)=$
$=\sum\limits_{(i,j)\in E}\biggl{(}\sum\limits_{c\in C(s,j)}(-1)^{\ell(c)+1}d^{\ell(c)+2}\biggl{)}=\sum\limits_{c\in C(s,i)}(-1)^{\ell(c)+1}d^{\ell(c)+2}\neq 0.$
So, the condition (5) is met for $i\!\in\mathcal{F}^{(k+1)}$. Let us show that $x_{i}^{(k+1)}\!=\!0$ if $i\!\not\in\mathcal{F}^{(k+1)}$ and the vertex $i$ was not visted at previous $k$ iterations of the combinatorial BFS. This means that among the vertices $j$ adjacent to $i$ there are no ones for which there have been simple chains of length $k$ connecting $s$ and $j$ traversed in the course of the previous $k$ iterations. Since Lemma 1 is true for $k$, this means that for all such $j$ we have $x_{j}^{(k)}\!=\!0$. Therefore,
@@
Thus, by (5) and by Lemma 1, the iterations (6) produce the same frontiers $\mathcal{F}^{(k)}$, $k\!=\!1,2,\ldots$, as iterations of the combinatorial BFS produce. This means that we obtain the same graph traversals during execution of the iterations. The Theorem 1 follows from that.
@@ 6 Gauss-Seidel iterations as implementation
6.1 Computing components of the state vector using traversals
of the chains outgoing from the current frontier’s verices
@@ 6.1 Computing components of the state vector using traversals
Considering the $(k+1)$-th iteration of the combinatorial CCS and iteration of the form (7) with the same number, we will use the following notations. Unlike to the previous section, we denote as $C(\tilde{s},i)$ the set of simple chains connecting the vertices $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ with vertices $i\!\in\!\mathcal{F}^{(k+1)}$. Along edges of these chains transitions are made at $(k+1)$-th iteration of the combinatorial CCS. Let $c+(j,i)$ denotes the chain obtained from $c$ by connecting its last vertex $j$ to vertex $i$ by the edge $(j,i)\!\in\!E$.
@@
All computations in iterations (6) and (7) can be decomposed into computations that implemented in accordance with traversals of individual chains (not necessary simple) connecting the starting vertex and the vertex $i$, for which the value $x_{i}^{(k+1)}$ is computing. We shall consider only simple chains since it will be enough for the proofs of the followed lemma and theorem. Indeed, when implemeting computations of the form (6) or (7), the computations for some set of walks from starting vertex to other vertices are constitute the set of all individual arithmetic operations implemented in the course of one iteration.
@@
Chains that can be traversed in the course of one iteration of combinatorial CCS, that is, chains $c\!=\!\{i_{0},i_{1},\ldots,i_{\ell(c)}\}\!\in\!C(\tilde{s},i)$, where $i_{0}\!=\!\tilde{s}$, are chains of the two types. A chain $c\!\in\!C(\tilde{s},i)$ is a chain of type (I) if $i_{0}\!<\!i_{1}$, and the chain type is (II) if $i_{0}\!>\!i_{1}$. For both type chains, maybe except the first edge $(i_{0},\!i_{1})$, we have $i_{j-1}\!<\!i_{j}$ for $j\!=\!\overline{1,\ell(c)}$. Transition along the edge $(i_{0},i_{1})$ is implemented at the step 3 of the combinatorial CCS and transitions along subsequent edges of the chain are performed at the step 4.
@@
For the vertices $\tilde{s},i\!\in\!V$, let $\ell(\tilde{s},i)\!=\!\max\{\ell(c)\,|\,c\in C(\tilde{s},i)\}$. $\ell(\tilde{s},i)$ is the maximum length of a simple chain outgoing from $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ along edges of which the vertex $i\!\in\!\mathcal{F}^{(k+1)}$ is reached at the $(k+1)$-th iteration.
@@
Let $i_{0}\!=\!\tilde{s}$, $i_{\ell(c)}\!=\!i$ in the chain $c\!=\!\{i_{0},\ldots,i_{\ell(c)}\}$ and $v_{\tilde{s}}$ is the value that is transmitted through the chain $c$ that connects the vertex $\tilde{s}$ and the vertex $i$. As a result of this transmission of $v_{\tilde{s}}$, by traversing the chain $c$, we obtain the contribution $x_{i,c}^{(k+1)}$ into the whole value of $x_{i}^{(k+1)}$ by this chain. There must be chains through which the non-zero value $v_{\tilde{s}}$ is transmitted to the vertex $i$ to have $x_{i}^{(k+1)}\!\neq\!0$.
@@
So, for $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ and $c\!\in C(\tilde{s},i)$ we define $x_{i,c}^{(k+1)}$ as the contribution of the value $v_{\tilde{s}}$ which transmitted along the chain $c\!\in C(\tilde{s},i)$ from $\tilde{s}$ to $i$:
@@
If $C(\tilde{s},i)\!=\!\varnothing$ then $x_{i(\tilde{s})}^{(k+1)}\!=\!0$ since there are no chains, through which the contributution $v_{\tilde{s}}$ can be transmitted to vertex $i$ from vertex $\tilde{s}$. The value $x_{i(\tilde{s})}^{(k+1)}$ is the sum of all contributions to the value of $x_{i}^{(k+1)}$ which are transmitted through all chains in $C(\tilde{s},i)$ at the course of the $(k+1)$-th iteration (7).
Note that during implementation of the iterartions (7), reaching vertices from $\mathcal{F}^{(k+1)}$ can be done not only along simple chains. Traversal of a chain of the type (II) may cause to correct chain that otgoes from some vertex $j$ already reached at the course of an iteration (7), at the same iteration will repeatedly pass through the vertex $\tilde{s}$ from which $j$ was reached. The value obtained along traversals of a such loops will be added to the initial value of $x_{\tilde{s}}^{(k)}$, further transmitted along the chains of $C(\tilde{s},i)$. That is, it will be included in $v_{\tilde{s}}$ from (8) while $\tilde{s}$ is revisited. This happens when $\tilde{s}\!>\!j$.
We show such situation on the Fig. 1.1.a) and Fig. 1.1.b). The equations (7) for these graph with starting vertex $2$ are the following:
$\left\{\begin{array}[]{l}x_{1}^{(k+1)}=x_{2}^{(k)}(-d),\\ x_{2}^{(k+1)}=\left(-1+x_{1}^{(k+1)}+x_{3}^{(k)}\right)(-d),\\ x_{3}^{(k+1)}=x_{2}^{(k+1)}(-d).\\ \end{array}\right.$
For the 1-st iteration, $x_{1}^{(0)}\!=\!0$, $x_{2}^{(0)}\!=\!d$, $x_{3}^{(0)}\!=\!0$, thus
$x_{3}^{(1)}=x_{2}^{(1)}(-d)=\left(-1+x_{1}^{(1)}\right)d^{2}=$
$=\left(-1+x_{2}^{(0)}(-d)\right)d^{2}=(-1-d^{2})d^{2}=-d^{2}-d^{4}.$
The contribution in $x_{3}^{(1)}$ of $v_{s}\!=\!d$ transmitted by the cycle $\{2,1,2\}$ gives us $-d^{4}$ for the whole walk $c_{1}\!=\!\{2,1,2,3\}$ at the $1$-st iteration. Adding it with the contribution in $x_{3}^{(1)}$ which is transmitted by the simple chain $c_{2}\!=\!\{1,3\}$ consisted of one edge $(1,3)$, we obtain the value of $x_{3}^{(1)}$ at the 1-st iteration as the sum of contributions in $x_{3}^{(1)}$ transmitted from the vertex $2$ by $3$ the walk $c_{1}$ and by the simple chain $c_{2}$:
$x_{3}^{(1)}=x_{3,c_{1}}^{(1)}+x_{3,c_{2}}^{(1)}=-d^{4}-d^{2}.$
Note that, after passing through such a loop (cycle), the contribution equally propagate over all simple chains outgoing from the vertex.
For the graph in Fig. 1.2, such loops (cycles) are absent in the case of the starting vertex $1$. The figure shows how the contributions added to the vertex $5$ along two correct chains that are traversed during one CSS iteration. These are the chains $c_{1}\!=\!\{1,2,5\}$ and $c_{2}\!=\!\{1,3,4,5\}$. We have
$x_{5}^{(1)}=x_{5,c_{1}}^{(1)}+x_{5,c_{2}}^{(1)}=d^{3}-d^{4}.$
@@ 6.2 Gauss-Seidel iterations as implementation
Lemma 2 states that obtaining the vector $x^{(k+1)}$ at the $(k+1)$-th iteration (7) is equivalent to obtaining it using all simple chains from $C(\tilde{s},i)$ in (9) for all $\tilde{s}\!\in\!\mathcal{F}^{(k)}$.
Proving the following lemma and theorem, we do not consider the situation of having cycles starting and ending at the vertex $\tilde{s}$, an example of which is given above. In this case, the contribution transmitted further after passing through such a loop (cycle) will equally transmitted over all correct simple chains outgoing from the vertex after passing the vertex again. This does not change the equality or inequality to zero of the sum of contributions transmitted along the chains in the course of (7). And therefore it does not change the equality or inequality to zero of $x_{i}^{(k+1)}$.
@@
$x_{i}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}x_{i(\tilde{s})}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}\sum\limits_{c\in C(\tilde{s},i)}x_{i,c}^{(k+1)}=\sum\limits_{\tilde{s}\in\mathcal{F}^{(k)}}\sum\limits_{c\in C(\tilde{s},i)}v_{\tilde{s}}\cdot(-d)^{\ell(c)},$ (11)11( 11 )
where $v_{\tilde{s}}\!=\!x_{\tilde{s}}^{(k)}$ is the value transmitted from the vertex $\tilde{s}\!\in\!\mathcal{F}^{(k)}$ by chains from $C(\tilde{s},i)$.
@@
Let us show that Lemma 2 is true for the $1$-st iteration, that is, for $k\!=\!1$. The frontier at the $1$-st iteration consists of a single starting vertex $s$: $\mathcal{F}^{(0)}\!=\!\{s\}$. Let $i\!\in\!\mathcal{F}^{(1)}$ be such that $\ell(s,i)\!=\!1$. Then $C(s,i)$ consists of a single chain $c\!=\!\{s,i\}$ of length $1$, i.e., it consists of the one edge $(s,i)\!\in\!E$. Thus we have $x_{i,c}^{(1)}\!=\!-d^{2}$ by (8), since $v_{s}\!=d$ at the $1$-st iteration (7).
For the $1$-st iteration (7), we have $x_{j}^{(0)}\!=\!0$ for all $j\!>\!i$, $(j,i)\!\in\!E$, $j\!\neq s$. We also have $x_{j}^{(1)}\!=\!0$ for $j\!<\!i$, $(j,i)\!\in\!E$, $j\!\neq\!s$. Let us assume that $x_{j}^{(1)}\!\neq\!0$ for such $j$. Consequently, there is a chain of the type (I) or type (II) connected vertices $s$ and $j$ such that, after substitutions of $x_{s}^{(0)}\!=\!d$ into equations (7) with numbers equal to numbers of vertices of such chain, we obtain $x_{j}^{(1)}\!\neq\!0$. But, in this case, we have $\ell(s,i)\!>\!1$ since $(j,i)\!\in\!E$ which contradicts assumption that $\ell(s,i)\!=\!1$.
So all the terms in (10) are zero except $x_{s}^{(0)}$, if $i\!<\!s$, or except $x_{s}^{(1)}$, if $i\!>\!s$, and we have
@@
where $\alpha\!=\!1$ if $(s,i)\!\in\!E$, $s\!>\!i$, and in this case we have $\{s,i\}\!\in\!C(s,i)\!\neq\!\varnothing$. If $\alpha\!=\!0$ and there are $x_{j}^{(0)}\!\neq\!0$, $j\!\neq\!s$, then there are other chains in $C(s,i)$. For all such $j$ we have $C(s,j)\!\neq\!\varnothing$, and, for $c\!\in\!C(s,j)$, $\ell(c)\!\leq\!\ell(s,j)\!\leq\!l$ since $\ell(s,i)\!=\!l+1$. Thus, by induction asssumption on $l$, we have:
@@
Since $x_{i,c}^{(1)}\!=\!x_{j,c^{\prime}}^{(1)}\cdot(-d)$ for $c^{\prime}\!\in C(s,j)$ such that $c\!=\!c^{\prime}+(j,i)$, $c\!\in C(s,i)$, then, substituting (13) into (12) and having $v_{s}\cdot(-d)\!=\!x_{i,c}^{(1)}$ for $c\!=\!\{s,i\}$, we obtain
@@
Consider $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!1$ for a vertex $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. For this case, $C(\tilde{s},i)$ consists of one chain that is the edge $(\tilde{s},i)$ which, in (10), corresponds to a non-zero term of the form
@@
Suppose that it is true for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!\leq l$ for $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. Let us show that it will also be true for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!l+1$. Carrying out induction on $l$ we show that for all vertices reached at the $(k+1)$-th iteration, the statement of the lemma is true.
@@
If $j\!<\!i$ and $x_{j}^{(k+1)}\!\neq\!0$, then $C(\tilde{s},j)\!\neq\!\varnothing$ since, by the induction assumption on $k$, only components for vertices reached at iterations up to the $k$-th can be nonzero in state vector, and to obtain a nonzero value $x_{j}^{(k+1)}$ in the right side of (10), correct chains connecting such vertices and $j$ are needed. The length of each chain in $C(\tilde{s},j)$ is less than or equal to $l$, since otherwise $\ell(\tilde{s},i)\!>\!l+1$.
Therefore, by induction on $l$, we have
@@
Consider the second term $S_{2}$ in the right-hand side of (10):
@@
Adding $S_{1}$ and $S_{2}$, we obtain from (15) and (16) that
@@
This means that Lemma 2 is true at the $(k+1)$-th iteration for all $i\!\in\!\mathcal{F}^{(k+1)}$ such that $\ell(\tilde{s},i)\!=\!l+1$. Thus, by induction on $l$, Lemma 2 is true for all vertices in $\mathcal{F}^{(k+1)}$, i.e., it is true for the $(k+1)$-th iteration. It means that, by the induction on $k$, it is true for an arbitrary $k$ ($k\!\leq\!n$).
@@
where $v_{\tilde{s}}\!=O(d^{\ell(s,\tilde{s})})\!\neq 0$, and where $\ell(s,\tilde{s})$ is the maximum length of a simple chain connecting the starting vertex $s$ and the vertex $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. By definition (8), when traversing along each edge of a such chain, the value $v_{\tilde{s}}$ which is a sum of powers $d$ taken with some coefficients, is multiplied by $-d$. In this case, the same powers of $d$ in the resulting values $x_{i,c}^{(k+1)}$ have the same sign, since they obtained by the same number of multiplications by $-d$ implemented by Chain traversal algorithm for chains of the same lengths. Therefore, if $C(\tilde{s},i)\!\neq\!\varnothing$, then $x_{i}^{(k+1)}\!\neq\!0$ and condition (5) is met for $i\!\in\!\mathcal{F}^{(k+1)}$ of the combinatorial CCS after the $(k+1)$-th iteration.
Let $i\!\in V\setminus\mathcal{C}^{(k+1)}$ for the combinatorial CCS, that is, the vertex $i$ is not reached after the $(k+1)$-th iteration. This means that $C(\tilde{s},i)\!=\!\varnothing$ for every $\tilde{s}\!\in\!\mathcal{F}^{(k)}$. By the proven Lemma 2, it follows from this that